A solution containg 4.50 g of a non electrolyte (i=1) dissolved in 125g of water freezes at -0.372°C. Calculate the molar mass of the solute.


Step 1: Firsy compute the molarity of the freezing point equation



Step 2: Then, from the definition of molality, compute the number of moles solute, n(solute), in the sample.



Step 3: Solve for the molar mass


1
Expert's answer
2021-04-29T07:22:09-0400

∆T = KF·Cm

where KF is the molal freezing point depression constant (for water KF = 1.86);

Cm is the molality of the solute.

Cm = ∆T / KF = 0.372 / 1.86 = 0.2 m

Cm = n / m

n = Cm x m = 0.2 x 0.125 = 0.025 mol

n = m / M

Msolute= m / n = 4.50 / 0.025 = 180 g/mol